OpenStudy (anonymous):

I have a question. (-5i) (-6-5i)

OpenStudy (anonymous):

DUDE SHUT UP

OpenStudy (anonymous):

30i+5i^2=-5+30i

OpenStudy (anonymous):

(-5i)*(-6-5i) = 30i + 25i^2 = 30i - 25 = -25 + 30i

OpenStudy (precal):

Sorry ndog33 Mandolino's solution is correct.